Run Coordination Meetings:Summer2022

From GlueXWiki
Jump to: navigation, search

Summer 2022 CPP/NPP run

Connect Information

Meeting are via ZoomGov: connect

Meeting ID: 160 199 0060 / Passcode 8394

Beam and Hall configuration

  • Experiments: NPP and CPP: June 8th- July 27th 2022. (Followed by PrimEx, start on Aug. 4th 20202).
  • Beam energy expected:
    • Expected: 11.57 GeV.
  • Solenoid:
    • Standard, 1350A
  • Pair Spectrometer: 6/9 = 0.67 of its nominal field (569.4 A = 0.67x849.9 A)
  • Microscope range: [0.301-0.514] of the beam energy or [4.57-6.01 GeV]
  • CW Beam:
    • 250 MHz frequency.
    • Current: 1-100 nA (exceptionally up to 350nA if needed).
    • Chopper slit:  ??To minimize bleedthrough, use Hall B slit whenever they are at 5-pass (full run period except first 4 days). Use Hall A slit the first 4 days (Hall A starts later)??
  • Position Slow Lock on or FFB on (FFB doesn't operate below 50nA).
  • Radiators:
    • Goniometer radiators:
      • Diamond JD70-103 (50 μm - 4.1·10-4 R.L., 7×7 mm²) - New crystal - Not expected to be great according to rocking curve.
      • Diamond JD70-107 (55 μm - 4.6·10-4 R.L., 7×7 mm²) - New crystal - Not expected to be great according to rocking curve.
      • (Diamond JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) - Spent: used for the second part of GlueX's Fall 2018 run and Fall 2019/Spring 2020/Fall 2021)
      • (Diamond JD70-106 (50 μm - 4.1·10-4 R.L., 7×7 mm²) - Aligned and used shortly during Tagger Accidental test in Spring 2020. Did not display sharp coherent peak.)
      • (Diamond JD70-109 (50 μm - 4.1·10-4 R.L., 7×7 mm²) - New crystal - Not expected to be good according to rocking curve.)
      • Amorphous 40 μm Al radiator (44.9·10-5 R.L.)
      • V-wire (for TAC run).
    • Amorphous
      • 1.64 µm Al (1.86·10-5 R.L., see study)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Tagger quadrupole on (-4.2 A).
  • Collimator hole: Default running: 3.4 mm.
  • Targets:
    • Lead target: 0.3mm, 5.4·10-2 R.L.
  • PS converter: 750 μm
  • Miscellaneous:
    • PS test area: Pb-W module tests (A. Somov)

General Information

This document describes the run plan for the Hall D Summer 2022 run.

  1. June 1st - 7th: Electron beam restoration.
  2. June 8th - July 27th: CPP/NPP run


  • Accelerator overall plan and priorities:
    • Deliver physics beam to 4 halls. Hall A starts later, on July 7th??. Hall A pass(es): TBD. Halls B and or C will always be at 5th pass (except for 4 days for Hall B), so Hall D beam will always be RF-separated and at 250 MHz.
  • Hall D plan
    • Experiments: CPP and NPP
    • Tentative commissioning plan
    • "TAC" runs (2 or more)
    • Energy calibration run: 5min every 2 days with CW beam, I=100nA
    • Straight track run. Tentatively between a TAC run and beam studies (to save on overhead): Ramp down the solenoid during beam study. Do the straight track run (6h if the beam delivery is good). Follow with a TAC run while ramping back up the solenoid.
    • No invasive non-CPP/NPP tests are planned.
  • Expected staffing and responsibilities:
    • The Run Coordinator oversees the run:
      • June 1st-June 7th: 7 days: Alexandre Deur (accelerator restoration)
      • June 8th-June 15th, 7 days: Alexandre Deur
      • June 15th-June 22nd, 7 days: Mark-Macrae Dalton
      • June 22nd-June 29th, 7 days: Naomi Jarvis
      • June 29th-July 6th, 7 days: Benedikt Zihlmann
      • July 6th-July 13th, 7 days: Ilya Larin
      • July 13th-July 20th, 7 days: Elton Smith
      • July 20th-July 27th, 7 days: Rory Miskimen


Runplan

  1. Verify electron beam quality and establish photon beam (assume 1 day).
    • Ion Chambers trip threshold must be set using diamond JD70-105 as well as the 4 Al. radiators (1 on goni, 3 on amorphous radiator ladder)
    • Tune electron beam with radiator retracted
      • Radiator (amorphous and diamond) must be retracted. Collimator should be in blocking position.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 ~30 mrad/hr (tagger area, gammas, on the ground below the goniometer).
          • RAD102_P3 ~ 0.3 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 ~ 0. mrad/hr (Hall. Neutrons).
    • Establish good photon beam
    • Make sure tagger quadrupole is on and at -4.2 A. (It can be turned on only once the Hall is in power/beam permit.)
    • Do not turn on the Active Collimator HV: Richard Jones will decided when to turn them on.
      • Insert 2*10-5 RL radiator. Call MCC and ask to mask the FSD and turn off beam each time a radiator is moved
        • Typical levels are (for 2×10-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 35 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= ~ 50 mrad/hr (tagger area, gammas, on the ground below the goniometer)
          • RAD102_P3= ~2 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Verify beam position and envelope (Hovanes).
      • Ask MCC to set beam position on 5C11b to x=-0.5 mm and y=+1.0 mm.
      • Ask MCC for 5C11 and 5C11B HARP. Then do our radiator HARP scan, see instruction. If it looks good, proceed to next step. Otherwise, inform MCC.
      • Insert profiler (no need for masking FSD when an object is inserted in the photon line), run with 10×10-4 radiator and 50 nA. Beam on the profiler should be at x=0.0±0.5cm; y=0.0±0.5cm. If not, inform MCC.
    • Radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 75 nA and 100 nA. There is need to run the DAQ or stop the beam between each steps: the data are recorded by EPICS. However, the beam need to be off when moving to a different radiator for the next steps.
      • Run 5 min with the 10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 75 nA and 100 nA
      • Run 3 min with the 3*10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 75 nA and 100 nA
    • Photon flux optimization and Active Collimator calibration. (20 + 20 mins). Call Hovanes.
      • re-Insert the 10-4 radiator (Call MCC). Use about 100 nA beam current. This will provide the Act. Col. nominal positions for its high gain values.
      • Determine maximum photon beam transmission by doing 2D scans using the collimator x/y motions (Hovanes. 1h)
      • Repeat procedure with 2*10-5 radiator (Call MCC) and 50 nA beam current. This will provide the Act. Col. nominal positions for its low gain values.
      • Record on the white board what is the beam position on 5C11b and Active Collimator for optimal photon transmission, both for low and high Act. Col. gain.
      • Until Hovanes determines to optimal 5C11b and Active Collimator position, keep profiler in and beam locked (MCC's duty) on the nominal 5C11b and x=0,y=0 profiler position. Once Active Collimator nominal beam position is known, retract profiler and lock the beam (MCC) on the active collimator.
  2. Detector and beamline checkout.
    • Any time before starting production (beam quality is not important): microscope HV scan (2h, R. Jones).
    • Establish L1 trigger, step 1: BCal trigger needs to be re-established before the ToF HV scans below. (Alex. Somov. 2h).
  3. ToF calibration (Beni, 6h). In parallel, CDC HV scans (Naomi). Use ??4.5×10-4 Al. radiator (on goni)??. Beam current 100nA
    • Establish L1 trigger, step 2: CPP/NPP trigger (Alex. Somov. ??h).
    • PS and TAGH HV scan. (Alex. Somov. 1h).
    • Establish DAQ (Sergei, 2h).
    • General detector calibration (Sean Dobbs. (703-887-1776) 1h). Use 1×10-4 Al. radiator. Beam current 100nA
      • Run 15 min (solid time) in Long Mode (Mode 10)
      • Run 15 min (solid time) in Production Mode (Mode 9).
  4. 50/55 μm diamond (Start on JD70-10?) alignement (20nA). 2 polarization configurations?? (Hovanes. 8h ). If the first diamond does not look good, Hovanes will align the second one. We need to know how much time CPP/NPP leadership wants to allow for diamond alignment.
  5. CPP/NPP tentative commissioning plan
  6. CPP/NPP run
    • Standard configuration: ~50 μm diamond, 27 nA, 3.4 mm hole, TPol on 750 μm TPol convertor. 2h runs and no more, as it may be hard to correct for calibration drifts. ??Cycle over 5 sets of runs: 0o (twice, or alternate??), then 90o (twice, or alternate??) then amorphous runs.??
      • Amorphous runs: 4.5×10-4 Al. radiator (on goni). 27 nA
    • Log in run information in the shift summary (make a run list). For each production run, log useful comments on DAQ comment window, see [1]
    • Tagger quadrupole on
    • Harp scans (30 min): Every 2 days?? May need to be done more often because the 3.5mm collimator will be used.
    • Energy calibration runs: 5min every 2 days with CW beam, I=100nA
    • Empty target run(s). How many?? 27 nA??: data taking duration?? 30min overhead to prepare the target (29min58s for rapid access/exit, 2s to flip the target switch).
  7. TAC runs. At least 2 runs.
    • a few nA, thinest Al. radiator and ComCal.
    • To be coordinated with MCC.
    • Measure bleedthrough. If acceptable, (~nA or less), proceed with TAC-ComCal run. Otherwise, explore possibility for invasive TAC run.
  8. ??Long mode run. Once a week (20 min).??
  9. Amorphous run. ??Every set of 5 diam. orientation runs. 1h Standard production current. Time to switch from pol. to unpol.: 10min. time to switch from unpol. to pol.: 30min.??

Hall D situation room (current run plan)

Hall D counting room white board



Accelerator status

Hot Checkout

To be completed by Tues. May. 31st

16% ready ; 0% checked ; 84% not ready

Readiness for Summer 2022 run

  • Solenoid (Eugene/Beni).
    • Everything is operational.
    • Most tests are done. Ramped to 800A on 05/17/2022. Ready
  • Surveys (Tim/Mark Stevens/Scott).
    • TAC
    • Luminosity Monitor.
    • Beam pipe installed.
    • Target. Done
    • Profiler. Not necessary as the new profiler will not be in.
    • 3.4mm collimator. Done
    • Test bed for Pb-W modules: no need for survey.
    • GEM/TRD detector: no need for survey.
  • DAQ (Sergei).
    • Everything is working with cosmic runs.
  • Trigger (Alex Somov).
    • Ready, except needs to check that ToF is properly in the trigger.
  • Beam line
    • Diamond status (Tim/Hovanes). Installed.
      • Broken LED light that illuminates the radiators has been replaced
    • Tagger hodoscope (A. Somov).
      • Need to set-up the HV and check the counters.
    • Tagger microscope (R. Jones/Beni).
      • Power Supply fixed.
      • Richard will conduct dark rate tests during the upcoming weeks.
    • Harp software development (Hovanes).
      • Needed for new 3-wire Harp that has been installed. On-line analysis software essentially done.
      • Need to be included in the commissioning plan.
    • Goniometer (Hovanes).
      • The camera failed. Cause unknown. Hovanes changed it (we had a spare).
    • Collimator Motion (Hovanes).
      • Tested yesterday. Working fine.
    • Active collimator detectors(R. Jones/A. Deur)
      • Has been moved to 3.4mm hole.
      • Total time (all included) to switch between two holes during run is 2h.
      • A.C. is Ready for the run.
    • Profilers (Hovanes)
      • The new profiler is not ready. We will continue to use the present profiler. (Not a critical issue.)
    • Luminosity monitor (aka Active Target) (Hovanes).
      • Installed and surveyed.
    • Halo monitors (Hovanes)
      • Installed.
      • The one in front of the target needs to be connected yet.
    • PS system (Alex Somov)
      • It would be good to fix one SiPM, if Chris Stanislav is available.
    • Polarimeter (M. Dugger/Beni/Fernando).
      • Mike Dugger is on site and working on the TPol. Noise of unknown origin is seen on the scope. Not clear if it is real. Need more investigations.
    • Pb Target (Tim)
    • Tim wrote the procedure for moving the target in and out.
      • Installed and (re-)surveyed.
      • Per Beni suggestion a flag (yellow tape) visible with the Counting Room camera and giving the target in/out status has been installed.
    • Slow control (Hovanes).
      • Everything that CPP requested is implemented. Beni checked and it works
      • Made new tagger NMR (to accommodate new accelerator gui). It seems to be working fine but needs to be tested with magnet on.
      • May need more work if the new profiler is installed. If not, next and final step is HCO.
    • RF (Sergei)
      • Everything checked fine, but could not check the CAEN crate.
      • Will be checked for all detectors next week.
    • Radiation monitors (A. Deur)
    • x-ray W shield (Tim/Lubomir).
      • Installed and ready.
  • Detector systems
    • ST (Beni)
      • Not used for this run.
    • FDC (Lubomir).
      • Cosmic data show that two cards are not giving signals. Issue is with the discriminators connected to the wires (the raw detector signals are fine). Still checking.
      • Beni did a new baseline calibration (also for the CDC). They will run this week end with it.
      • FDC chiller was down. Lubomir restarted it.
    • CDC (Beni).
      • Took cosmic runs. Everything looks good with the CDC.
    • Gas system (Beni/Lubomir)
      • It is still suspected that the 1st FDC chamber is clogged. Lubomir is watching its efficiency.
      • We will use the gas analyzer during this run.
      • The 4He bottle was empty again. Josh to change the regulator.
    • BCal (Mark Dalton)
      • On and functioning (currently used for the cosmic run trigger).
      • Need to watch occupancy for ADS failures.
    • FCal (Mark Dalton)
      • On. Changed 25 low-gain PMTs.
    • ComCal (Alex Somov).
      • Not used for this run (CPP/NPP), but will be mounted for next one (PrimEx).
    • TAC (Alex Somov).
      • Will use 3×3 Pb-W blocks as new TAC, with signals from ComCal cables.
      • Beni will test signals to make sure the cables are plugged in correctly.
      • Installed and surveyed. Need to be cabled.
    • ToF (Beni).
      • 4 discriminators are not counting because cables were not connected but are now fixed.
    • MWPC (Muon detector) (Beni)
      • Up and running. All channels have signals. 1 channel is hot (as well as its 2 neighbors due to cross-talk). This was dealt with by increasing its threshold and the occupancy is now more or less flat.
    • Pb absorber (Elton)
      • Mounted.
      • Tim wr0te the procedure on how to change its thickness.
      • Need training to modify it. Tentatively given (M. Dalton) on June 3rd, 9:30am.
    • DIRC (Justin).
      • Not used for this run.
    • GEM/TRD (Sergei/Lubomir).
      • There will be probably no tests (even parasitic) done during the CPP run.
    • Monitoring (Alexander Austregesilo)
      • The Translation Table is in. Two new monitoring plots have been made available.
      • Took cosmics and the plots look fine.
      • High rate test scheduled for tomorrow (May 20th) using simulated data.
  • RSAD and ESAD Safety documents (Simon):
    • All documents ready.
    • OSP submitted and approved.
  • Phones

Live Camera Feeds from the Four Counting Houses

  1. These feeds SHOULD be open and visible at all times from one of the monitors in each of the counting houses.
  2. Management cautions AGAINST opening them from other locations or offsite, out of concern for available bandwidth.

Running Shifts Remote